OpenStudy (anonymous):

The function f(x) = 4.50x + 19 relates how much Sean pays to his scout club, f(x), to the number of patches, x, he earns in a month. What is the value and meaning of f(12)?

OpenStudy (jennithemeani):

f(12)=4.50(12)+19 f(12)=54+19 f(12)=73

OpenStudy (clalgee):

Please do not give direct answers :) Welcome to Openstudy @princesa_gemma !